# Solving General-Utility Markov Decision Processes in the Single-Trial Regime with Online Planning

### v1

@@ 4.1 Monte-Carlo tree search for GUMDPs in the single-trial regime
We employ an MCTS algorithm to solve the occupancy MDP. As described in Sec. 2.2, the search tree of the online planning algorithm comprises decision and chance nodes. In the context of the occupancy MDP, each decision node corresponds to an action $a\in\mathcal{A}$, while each chance node corresponds to a given state $\{s,{\boldsymbol{o}}\}\in\mathcal{S}\times\mathcal{O}$ of the occupancy MDP. At timestep $t$ of the interaction, the MCTS algorithm builds a planning tree rooted at the current state $\{s_{t},{\boldsymbol{o}}_{t}\}$, following the four phases outlined in Sec. 2.2 at each iteration.
Remark 2.
Assume that the objective function is bounded in its domain, i.e., $f_{\text{min}}\leq f({\boldsymbol{d}})\leq f_{\text{max}}$, for any ${\boldsymbol{d}}\in\Delta(\mathcal{S}\times\mathcal{A})$. Then, for any horizon $H\in\mathbb{N}$, the MCTS algorithm provably solves the occupancy MDP as the number of iterations of the algorithm per timestep grows to infinity. Furthermore, for any horizon $H\in\mathbb{N}$, we have that $\mathcal{R}(\pi_{\text{MCTS}})\leq 8\frac{L}{f_{\text{max}}-f_{\text{min}}}\gamma^{H}$ as the number of iterations of the algorithm per timestep grows to infinity, where we let $\pi_{\text{MCTS}}$ be the policy induced by the MCTS algorithm at each timestep. This result follows from our Theo. 2 and Theo. 6 in Kocsis and Szepesvári (2006) by rescaling the objective function to lie in the $[0,1]$ interval.
@@ 5 Experimental Results
In this section, we empirically assess the performance of the proposed MCTS-based algorithm for solving GUMDPS in the single-trial setting. Below, we provide a brief description of the considered tasks and environments. We refer to Appendix C for a complete description of our environments, experimental details, and complete experimental results.
@@
Tasks, environments, baselines, and experimental methodology
We consider three tasks: (i) maximum state entropy exploration (Hazan et al., 2019); imitation learning (Abbeel and Ng, 2004); and (iii) adversarial MDPs (Rosenberg and Mansour, 2019). We consider two sets of environments. The first set consists of the illustrative GUMDPs depicted in Fig. 1, each associated with one of the tasks. The second set of environments come from the OpenAI Gym library (Brockman et al., 2016). We consider the FrozenLake (FL), Taxi, and mountaincar (MC) environments. The framework of GUMDPs is defined over discrete state spaces; hence, we discretized the MC environment using a $10\times 10$ grid with equally-spaced bins. For the FL, Taxi, and MC environments, the task of imitation learning consists in imitating an approximately optimal policy. We let $\gamma=0.9$ and set $H=100$ for the illustrative GUMDPs and $H=200$ for the other environments. We perform 10 runs per experimental setting. We consider two baselines: (i) a random policy, $\pi_{\text{Random}}$; and (ii) the optimal policy for the infinite trials formulation (2), $\pi_{\text{Solver}}^{*}$, calculated by solving a constrained optimization problem with objective $f$ via a standard optimization solver. We denote the policy induced by our MCTS algorithm as $\pi_{\text{MCTS}}$ and consider 4000 iterations per timestep (results with other numbers of iterations in the Appendix C).
Experimental results discussion
@@ 6 Conclusion & Limitations
In this work, we contribute with the first approach to solve infinite-horizon discounted GUMDPs in the single-trial regime. In Sec. 3, we provided the fundamental results underpinning policy optimization in the discounted single-trial regime. Then, in Secs. 4 and 5, we explored how we can resort to online planning techniques, in particular MCTS, to solve discounted GUMDPs in the single-trial regime. Our work takes a first step towards a broader application of GUMDPs in real-world settings where the agent’s performance is typically evaluated under a single trial.
The key limitations of our approach to solve GUMDPs in the single-trial regime are: (i) the MCTS algorithm requires a simulator of the environment to sample transitions; and (ii) the size of the matrix that keeps track of the running occupancy may be impractical for GUMDPs with large state and action spaces. We believe such limitations should be addressed by future work, for example, by investigating methods to compress the running occupancy (e.g., by employing state aggregation). We also believe extending our analysis to environments with continuous state spaces is interesting.
@@ Appendix A Lipschitz constants
Table 2: Common objective functions found in the GUMDPs literature. In ()($\dagger$)( † ) we assume ${\boldsymbol{d}}$ is lower bounded by $\epsilon$ satisfying $0<\epsilon<e^{-2}$.
Task Objective ($f({\boldsymbol{d}})$) Lipschitz constant ($L$)
MDPs/RL ${\boldsymbol{d}}^{\top}{\boldsymbol{c}},\quad{\boldsymbol{c}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ $\max_{s,a}|c(s,a)|$
Pure exploration ${\boldsymbol{d}}^{\top}\log({\boldsymbol{d}})$ $|\log(\epsilon)+1|\;(\dagger)$
Imitation learning $\|{\boldsymbol{d}}-{\boldsymbol{d}}_{\beta}\|_{2}^{2},\quad{\boldsymbol{d}}_{\beta}\in\Delta(\mathcal{S}\times\mathcal{A})$ $4$
Adversarial MDPs $\max_{k\in\{1,\ldots,K\}}{\boldsymbol{d}}^{\top}{\boldsymbol{c}}_{k}$ $\max_{k\in\{1,\ldots,K\}}\left\{\max_{s,a}|c_{k}(s,a)|\right\}$
Objective function $f({\boldsymbol{d}})={\boldsymbol{c}}^{\top}{\boldsymbol{d}}$
@@
$|f({\boldsymbol{d}}_{1})-f({\boldsymbol{d}}_{2})|=|{\boldsymbol{c}}^{\top}({\boldsymbol{d}}_{1}-{\boldsymbol{d}}_{2})|=\sum_{s,a}|c(s,a)||d_{1}(s,a)-d_{2}(s,a)|\leq\max_{s,a}|c(s,a)|\|{\boldsymbol{d}}_{1}-{\boldsymbol{d}}_{2}\|_{1}.$
Objective function $f({\boldsymbol{d}})={\boldsymbol{d}}^{\top}\log({\boldsymbol{d}})$
We assume ${\boldsymbol{d}}$ is lower bounded by $\epsilon$, i.e., $d(s,a)\geq\epsilon$ with $0<\epsilon<e^{-2}$ for all $s\in\mathcal{S},a\in\mathcal{A}$. We let $f({\boldsymbol{d}})=\sum_{s,a}g(d(s,a))$, for $g(x)=x\log(x)$. We note that, $g^{\prime}(x)=\log(x)+1$ and it holds for any $x\in[\epsilon,1]$ that $|g^{\prime}(x)|\leq|\log(\epsilon)+1|$. Thus, for any $x_{1},x_{2}\in[\epsilon,1]$ we have that
[…]
@@
Thus, for any ${\boldsymbol{d}}_{1},{\boldsymbol{d}}_{2}\in\Delta(\mathcal{S}\times\mathcal{A})$ lower bounded by $0<\epsilon<e^{-2}$, it holds that
$|f({\boldsymbol{d}}_{1})-f({\boldsymbol{d}}_{1})|$ $=\left|\sum_{s,a}\left(g(d_{1}(s,a))-g(d_{2}(s,a))\right)\right|$
@@
$=\left|\log(\epsilon)+1\right|\|{\boldsymbol{d}}_{1}-{\boldsymbol{d}}_{2}\|_{1}$
@@
Objective function $f({\boldsymbol{d}})=\|{\boldsymbol{d}}-{\boldsymbol{d}}_{\beta}\|_{2}^{2}$
It holds that $\nabla f({\boldsymbol{d}})=2({\boldsymbol{d}}-{\boldsymbol{d}}_{\beta})$. Now,
$\max_{{\boldsymbol{d}}\in\Delta(\mathcal{S}\times\mathcal{A})}\|\nabla f({\boldsymbol{d}})\|_{1}=2\max_{{\boldsymbol{d}}\in\Delta(\mathcal{S}\times\mathcal{A})}\|{\boldsymbol{d}}-{\boldsymbol{d}}_{\beta}\|_{1}\leq 2\max_{{\boldsymbol{d}}_{1},{\boldsymbol{d}}_{2}\in\Delta(\mathcal{S}\times\mathcal{A})}\|{\boldsymbol{d}}_{1}-{\boldsymbol{d}}_{2}\|_{1}=4.$
@@ B.1 Proof of Theorem 1
To prove our result, we consider the GUMDP depicted in Fig. 2. To simplify our proof, we consider state-dependent occupancies and denote an occupancy for the GUMDP above with the vector ${\boldsymbol{d}}=[d(s^{0}),d(s^{1}),d(s^{2})]$. The objective function is $f({\boldsymbol{d}})={\boldsymbol{d}}^{\top}{\boldsymbol{A}}{\boldsymbol{d}}$, where $A=\text{diag}([0,1,1])$, which is Lipschitz (over $\Delta(\mathcal{S})$) and convex. Under any trajectory it holds that $d(s^{0})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{2t+1}=\frac{(1-\gamma)\gamma}{1-\gamma^{2}}$. Hence, it holds that, under any trajectory, $d(s^{1})+d(s^{2})=1-\frac{(1-\gamma)\gamma}{1-\gamma^{2}}=\frac{1-\gamma}{1-\gamma^{2}}$. Thus, we focus our attention to the value of the occupancy at state $s^{1}$, $d(s^{1})$, and let $d(s^{2})=\frac{1-\gamma}{1-\gamma^{2}}-d(s^{1})$. With this, we can define our objective as a function of $d(s^{1})$ only by letting $f(d(s^{1}))=d(s^{1})^{2}+(\frac{1-\gamma}{1-\gamma^{2}}-d(s^{1}))^{2}$.
@@
=$\overset{(a)}{=}\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{S}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg{|}\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{1}),\\ \ldots,\\ \mathrm{s}_{3}\sim P^{\mathrm{a}_{2}}(\cdot|s_{2}),\mathrm{a}_{3}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{3}),\ldots\end{subarray}\right]$
=$\overset{(b)}{=}\mathbb{E}\left[\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{S}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg{|}\begin{subarray}{c}\mathrm{a}_{5}\sim\pi_{\text{S}}(\cdot|s^{0}),\\ \mathrm{s}_{6}\sim P^{\mathrm{a}_{5}}(\cdot|s^{0}),\ldots\end{subarray}\right]\Bigg{|}\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{1}),\\ \ldots,\\ \mathrm{s}_{4}\sim P^{\mathrm{a}_{3}}(\cdot|s_{3}),\mathrm{a}_{4}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{4})\end{subarray}\right]$
=$\overset{(c)}{=}\mathbb{E}\left[\mathbb{E}\left[f\left((1-\gamma)\sum_{t=0}^{4}\gamma^{t}\mathbf{1}(\mathrm{s}_{t}=s^{1})+\gamma^{5}\tilde{\mathrm{d}}^{\pi_{\text{S}}}_{\mathrm{a}_{5},\mathrm{s}_{6},\mathrm{a}_{6},\ldots}(s^{1})\right)\Bigg{|}\begin{subarray}{c}\mathrm{a}_{5}\sim\pi_{\text{S}}(\cdot|s^{0}),\\ \mathrm{s}_{6}\sim P^{\mathrm{a}_{5}}(\cdot|s^{0}),\ldots\end{subarray}\right]\Bigg{|}\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{1}),\\ \ldots,\\ \mathrm{s}_{4}\sim P^{\mathrm{a}_{3}}(\cdot|s_{3}),\mathrm{a}_{4}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{4})\end{subarray}\right]$
@@
=$\overset{(e)}{=}\mathbb{E}\Bigg{[}f\Bigg{(}(1-\gamma)\left(\mathbf{1}(\mathrm{s}_{0}=s^{1})+\gamma^{2}\mathbf{1}(\mathrm{s}_{2}=s^{1})+\gamma^{4}\mathbf{1}(\mathrm{s}_{4}=s^{1})\right)$
@@
where in (a) we emphasized that the random vector $\mathbf{d}^{\pi_{\text{S}}}$ depends on random variables $\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots$; (b) follows from the fact that $\mathrm{s}_{5}=s^{0}$ with probability one for trajectories drawn from the GUMDP, i.e., for any trajectory the state at timestep $5$ is always $s^{0}$, and thus we can split and simplify the expectation. In step (c), we let $\tilde{\mathbf{d}}$ be the random vector defined as in (3) for the GUMDP depicted in Fig. 2, but where $p_{0}(s^{0})=1$ and zero otherwise. Step (d) follows from Jensen’s inequality since, for any values $(s_{0},a_{0},s_{1},a_{1},\ldots,s_{4},a_{4})\in(\mathcal{S}\times\mathcal{A})^{5}$ the random variables of the outer expectation can take, it holds that $\mathbb{E}[g(\tilde{\mathbf{d}}^{\pi_{\text{S}}})]\geq g(\mathbb{E}[\tilde{\mathbf{d}}^{\pi_{\text{S}}}])$ where we let $g(x)=f((1-\gamma)\sum_{t=0}^{4}\gamma^{t}\mathbf{1}(s_{t}=s^{1})+\gamma^{5}x)$, which is convex. Finally, step (e) follows from the fact that
@@
=$\overset{(a)}{=}\frac{1}{2}f\left((1-\gamma)\left(1+\frac{\gamma^{4}}{1-\gamma^{4}}\right)\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{4}}{1-\gamma^{4}}\right)$
@@
which can be verified using a software for symbolic/algebraic computation such as Mathematica [Wolfram Research, ]. We provide a snippet of the Mathematica code we used below.

### v2

@@ 4.1 Monte-Carlo tree search for GUMDPs in the single-trial regime
We employ an MCTS algorithm to solve the occupancy MDP. As described in Sec. 2.2, the search tree of the online planning algorithm comprises decision and chance nodes. In the context of the occupancy MDP, each decision node corresponds to an action $a\in\mathcal{A}$, while each chance node corresponds to a given state $\{s,{\bm{o}}\}\in\mathcal{S}\times\mathcal{O}$ of the occupancy MDP. At timestep $t$ of the interaction, the MCTS algorithm builds a planning tree rooted at the current state $\{s_{t},{\bm{o}}_{t}\}$, following the four phases outlined in Sec. 2.2 at each iteration. We provide the following two remarks, which further characterize the convergence and regret of our proposed MCTS-based approach for solving GUMDPs in the single-trial regime. To derive the two results below, we make the following assumption.
Assumption 2.
The objective function satisfies $f_{\text{min}}\leq f({\bm{d}})\leq f_{\text{max}}$, for any ${\bm{d}}\in\Delta(\mathcal{S}\times\mathcal{A})$.
Remark 2 (Convergence).
Under Assumption 2, for any horizon $H\in\mathbb{N}$, the MCTS algorithm provably solves the occupancy MDP as the number of iterations of the algorithm per timestep grows to infinity. More precisely, for any horizon $H\in\mathbb{N}$, we have that $\mathrm{OptGap}_{H}(\pi_{\text{MCTS}})=0$ as the number of iterations of the algorithm per timestep grows to infinity, where we let $\pi_{\text{MCTS}}$ be the policy induced by the MCTS algorithm at each timestep. This result follows from our Theo. 2 and Theo. 6 in Kocsis and Szepesvári (2006) by rescaling the objective function to lie in the $[0,1]$ interval. Thus, from Prop. 1, $\mathrm{OptGap}(\pi_{\text{MCTS}})\leq 8\frac{L}{f_{\text{max}}-f_{\text{min}}}\gamma^{H}$.
Remark 3 (Regret bound).
For a given number of episodes $n$ and a sequence of policies $(\pi_{1},\ldots,\pi_{n})$, let $\mathcal{R}(\pi_{1},\ldots,\pi_{n})=\frac{1}{n}\sum_{i=1}^{n}\textrm{OptGap}(\pi_{i})$ be the expected regret for the sequence of policies and $\mathcal{R}_{H}(\pi_{1},\ldots,\pi_{n})=\frac{1}{n}\sum_{i=1}^{n}\textrm{OptGap}_{H}(\pi_{i})$ the truncated expected regret for the sequence of policies. Under Assumption 2, for any horizon $H$, there exists an MCTS-based algorithm such that, when applied to solve the occupancy MDP with $n$ expansion steps, yields a sequence of policies $(\pi_{1},\ldots,\pi_{n})$ satisfying $\mathcal{R}_{H}(\pi_{1},\ldots,\pi_{n})\leq O(1/\sqrt{n})$. Furthermore, from Prop. 1, $\mathcal{R}(\pi_{1},\ldots,\pi_{n})\leq O(1/\sqrt{n}+\gamma^{H})$. This result follows from our Theo. 2 and the works of Shah et al. (2020) and Cömer et al. (2025), particularly Theo. 1 in Shah et al. (2020) and Theo. 1 in Cömer et al. (2025), which establish polynomial regret bounds for MCTS-based algorithms.
@@ 5 Experimental Results
In this section, we empirically assess the performance of the proposed MCTS-based algorithm for solving GUMDPs in the single-trial setting. Below, we provide a brief description of the considered tasks and environments. We refer to Appendix C for a complete description of our experiments.
Tasks, environments, baselines, and experimental methodology
We consider three tasks: (i) maximum state entropy exploration (Hazan et al., 2019); imitation learning (Abbeel and Ng, 2004); and (iii) adversarial MDPs (Rosenberg and Mansour, 2019). We consider two sets of environments. The first set consists of the illustrative GUMDPs depicted in Fig. 1, each associated with one of the tasks. The second set of environments come from the OpenAI Gym library (Brockman et al., 2016). We consider the FrozenLake (FL), Taxi, and mountaincar (MC) environments. The framework of GUMDPs is defined over discrete state spaces; hence, we discretized the MC environment using a $10\times 10$ grid with equally-spaced bins. For the FL, Taxi, and MC environments, the task of imitation learning consists in imitating an approximately optimal policy. We let $\gamma=0.9$ and set $H=100$ for the illustrative GUMDPs and $H=200$ for the other environments. We perform 10 runs per experimental setting. We consider two baselines: (i) a random policy, $\pi_{\text{Random}}$; and (ii) the optimal policy for the infinite trials formulation equation 2, $\pi_{\text{Solver}}^{*}$, calculated by solving a constrained optimization problem with objective $f$ via a standard optimization solver. We denote the policy induced by our MCTS algorithm as $\pi_{\text{MCTS}}$ and consider 4000 iterations per timestep (results with other numbers of iterations in the Appendix C). Our code can be found here
Experimental results discussion
@@ 6 Conclusion & Limitations
In this work, we contribute with the first approach to solve infinite-horizon discounted GUMDPs in the single-trial regime. In Sec. 3, we provided the fundamental results underpinning policy optimization in the discounted single-trial regime. Then, in Secs. 4 and 5, we explored how we can resort to online planning techniques, in particular MCTS, to solve discounted GUMDPs in the single-trial regime. Our work takes a first step towards a broader application of GUMDPs in real-world settings where the agent’s performance is typically evaluated under a single trial. The key limitations of our approach to solve GUMDPs in the single-trial regime are: (i) the MCTS algorithm requires a simulator of the environment to sample transitions; and (ii) the size of the matrix that keeps track of the running occupancy may be impractical for GUMDPs with large state and action spaces. We believe such limitations should be addressed by future work, for example, by investigating methods to compress the running occupancy.
Acknowledgments
This work was supported by Portuguese national funds through the Portuguese Fundação para a Ciência e a Tecnologia (FCT) under projects UID/50021/2025 and UID/PRR/50021/2025 (INESC-ID multi-annual funding), as well as AI-PackBot (project number 14935, LISBOA2030-FEDER-00854700). Pedro P. Santos acknowledges the FCT PhD grant 2021.04684.BD. Alberto Sardinha acknowledges the CNPq Research Productivity Fellowship (PQ), with reference 312699/2025-5. The authors also thank the lab managers at GAIPS for the support provided when running the computational experiments of this work.
@@ Appendix A Lipschitz constants
Table 2: Common objective functions found in the GUMDPs literature. In $(\dagger)$ we assume $\bm{d}$} is lower bounded by $\epsilon$ satisfying $0<\epsilon<e^{-2}$.
Task Objective ($f({\bm{d}})$) Lipschitz constant ($L$)
MDPs/RL ${\bm{d}}^{\top}{\bm{c}},\quad{\bm{c}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ $\max_{s,a}|c(s,a)|$
Pure exploration ${\bm{d}}^{\top}\log({\bm{d}})$ $|\log(\epsilon)+1|\;(\dagger)$
Imitation learning $\|{\bm{d}}-{\bm{d}}_{\beta}\|_{2}^{2},\quad{\bm{d}}_{\beta}\in\Delta(\mathcal{S}\times\mathcal{A})$ 44
Adversarial MDPs $\max_{k\in\{1,\ldots,K\}}{\bm{d}}^{\top}{\bm{c}}_{k}$ $\max_{k\in\{1,\ldots,K\}}\left\{\max_{s,a}|c_{k}(s,a)|\right\}$
Objective function $f({\bm{d}})={\bm{c}}^{\top}{\bm{d}}$
@@
$|f({\bm{d}}_{1})-f({\bm{d}}_{2})|=|{\bm{c}}^{\top}({\bm{d}}_{1}-{\bm{d}}_{2})|=\sum_{s,a}|c(s,a)||d_{1}(s,a)-d_{2}(s,a)|\leq\max_{s,a}|c(s,a)|\|{\bm{d}}_{1}-{\bm{d}}_{2}\|_{1}.$
Objective function $f({\bm{d}})={\bm{d}}^{\top}\log({\bm{d}})$
We assume $\bm{d}$} is lower bounded by $\epsilon$, i.e., $d(s,a)\geq\epsilon$ with $0<\epsilon<e^{-2}$ for all $s\in\mathcal{S},a\in\mathcal{A}$. We let $f({\bm{d}})=\sum_{s,a}g(d(s,a))$, for $g(x)=x\log(x)$. We note that, $g^{\prime}(x)=\log(x)+1$ and it holds for any $x\in[\epsilon,1]$ that $|g^{\prime}(x)|\leq|\log(\epsilon)+1|$. Thus, for any $x_{1},x_{2}\in[\epsilon,1]$ we have that
$|g(x_{1})-g(x_{2})|$ $=\left|\int_{x_{2}}^{x_{1}}g^{\prime}(x)dx\right|=\left|\int_{\min\{x_{1},x_{2}\}}^{\max\{x_{1},x_{2}\}}g^{\prime}(x)dx\right|$
$\leq\int_{\min\{x_{1},x_{2}\}}^{\max\{x_{1},x_{2}\}}\left|g^{\prime}(x)\right|dx\leq\int_{\min\{x_{1},x_{2}\}}^{\max\{x_{1},x_{2}\}}\left|\log(\epsilon)+1\right|dx$
@@
Thus, for any ${\bm{d}}_{1},{\bm{d}}_{2}\in\Delta(\mathcal{S}\times\mathcal{A})$ lower bounded by $0<\epsilon<e^{-2}$, it holds that
$|f({\bm{d}}_{1})-f({\bm{d}}_{1})|$ $=\left|\sum_{s,a}\left(g(d_{1}(s,a))-g(d_{2}(s,a))\right)\right|$
@@
$=\left|\log(\epsilon)+1\right|\|{\bm{d}}_{1}-{\bm{d}}_{2}\|_{1}$
@@
Objective function $f({\bm{d}})=\|{\bm{d}}-{\bm{d}}_{\beta}\|_{2}^{2}$
It holds that $\nabla f({\bm{d}})=2({\bm{d}}-{\bm{d}}_{\beta})$. Now,
$\max_{{\bm{d}}\in\Delta(\mathcal{S}\times\mathcal{A})}\|\nabla f({\bm{d}})\|_{1}=2\max_{{\bm{d}}\in\Delta(\mathcal{S}\times\mathcal{A})}\|{\bm{d}}-{\bm{d}}_{\beta}\|_{1}\leq 2\max_{{\bm{d}}_{1},{\bm{d}}_{2}\in\Delta(\mathcal{S}\times\mathcal{A})}\|{\bm{d}}_{1}-{\bm{d}}_{2}\|_{1}=4.$
@@ B.1 Proof of Theorem 1
To prove our result, we consider the GUMDP depicted in Fig. 2. To simplify our proof, we consider state-dependent occupancies and denote an occupancy for the GUMDP above with the vector ${\bm{d}}=[d(s^{0}),d(s^{1}),d(s^{2})]$. The objective function is $f({\bm{d}})={\bm{d}}^{\top}{\bm{A}}{\bm{d}}$, where $A=\text{diag}([0,1,1])$, which is Lipschitz (over $\Delta(\mathcal{S})$) and convex. Under any trajectory it holds that $d(s^{0})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{2t+1}=\frac{(1-\gamma)\gamma}{1-\gamma^{2}}$. Hence, it holds that, under any trajectory, $d(s^{1})+d(s^{2})=1-\frac{(1-\gamma)\gamma}{1-\gamma^{2}}=\frac{1-\gamma}{1-\gamma^{2}}$. Thus, we focus our attention to the value of the occupancy at state $s^{1}$, $d(s^{1})$, and let $d(s^{2})=\frac{1-\gamma}{1-\gamma^{2}}-d(s^{1})$. With this, we can define our objective as a function of $d(s^{1})$ only by letting $f(d(s^{1}))=d(s^{1})^{2}+(\frac{1-\gamma}{1-\gamma^{2}}-d(s^{1}))^{2}$.
@@
$\overset{(a)}{=}\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{S}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg|\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{1}),\\ \ldots,\\ \mathrm{s}_{3}\sim P^{\mathrm{a}_{2}}(\cdot|s_{2}),\mathrm{a}_{3}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{3}),\ldots\end{subarray}\right]$
$\overset{(b)}{=}\mathbb{E}\left[\mathbb{E}\left[f(\mathrm{d}^{\pi_{\text{S}}}_{\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots}(s^{1}))\Bigg|\begin{subarray}{c}\mathrm{a}_{5}\sim\pi_{\text{S}}(\cdot|s^{0}),\\ \mathrm{s}_{6}\sim P^{\mathrm{a}_{5}}(\cdot|s^{0}),\ldots\end{subarray}\right]\Bigg|\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{1}),\\ \ldots,\\ \mathrm{s}_{4}\sim P^{\mathrm{a}_{3}}(\cdot|s_{3}),\mathrm{a}_{4}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{4})\end{subarray}\right]$
$\overset{(c)}{=}\mathbb{E}\left[\mathbb{E}\left[f\left((1-\gamma)\sum_{t=0}^{4}\gamma^{t}\mathbf{1}(\mathrm{s}_{t}=s^{1})+\gamma^{5}\tilde{\mathrm{d}}^{\pi_{\text{S}}}_{\mathrm{a}_{5},\mathrm{s}_{6},\mathrm{a}_{6},\ldots}(s^{1})\right)\Bigg|\begin{subarray}{c}\mathrm{a}_{5}\sim\pi_{\text{S}}(\cdot|s^{0}),\\ \mathrm{s}_{6}\sim P^{\mathrm{a}_{5}}(\cdot|s^{0}),\ldots\end{subarray}\right]\Bigg|\begin{subarray}{c}\mathrm{s}_{0}\sim p_{0},\mathrm{a}_{0}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{0}),\\ \mathrm{s}_{1}\sim P^{\mathrm{a}_{0}}(\cdot|s_{0}),\mathrm{a}_{1}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{1}),\\ \ldots,\\ \mathrm{s}_{4}\sim P^{\mathrm{a}_{3}}(\cdot|s_{3}),\mathrm{a}_{4}\sim\pi_{\text{S}}(\cdot|\mathrm{s}_{4})\end{subarray}\right]$
@@
$\overset{(e)}{=}\mathbb{E}\Bigg[f\Bigg((1-\gamma)\left(\mathbf{1}(\mathrm{s}_{0}=s^{1})+\gamma^{2}\mathbf{1}(\mathrm{s}_{2}=s^{1})+\gamma^{4}\mathbf{1}(\mathrm{s}_{4}=s^{1})\right)$
@@
where in (a) we emphasized that the random vector $\mathbf{d}^{\pi_{\text{S}}}$ depends on random variables $\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots$; (b) follows from the fact that $\mathrm{s}_{5}=s^{0}$ with probability one for trajectories drawn from the GUMDP, i.e., for any trajectory the state at timestep $5$ is always $s^{0}$, and thus we can split and simplify the expectation. In step (c), we let $\tilde{\mathbf{d}}$ be the random vector defined as in equation 3 for the GUMDP depicted in Fig. 2, but where $p_{0}(s^{0})=1$ and zero otherwise. Step (d) follows from Jensen’s inequality since, for any values $(s_{0},a_{0},s_{1},a_{1},\ldots,s_{4},a_{4})\in(\mathcal{S}\times\mathcal{A})^{5}$ the random variables of the outer expectation can take, it holds that $\mathbb{E}[g(\tilde{\mathbf{d}}^{\pi_{\text{S}}})]\geq g(\mathbb{E}[\tilde{\mathbf{d}}^{\pi_{\text{S}}}])$ where we let $g(x)=f((1-\gamma)\sum_{t=0}^{4}\gamma^{t}\mathbf{1}(s_{t}=s^{1})+\gamma^{5}x)$, which is convex. Finally, step (e) follows from the fact that
@@
$\overset{(a)}{=}\frac{1}{2}f\left((1-\gamma)\left(1+\frac{\gamma^{4}}{1-\gamma^{4}}\right)\right)+\frac{1}{2}f\left(\frac{(1-\gamma)\gamma^{4}}{1-\gamma^{4}}\right)$
@@
which can be verified using a software for symbolic/algebraic computation such as Mathematica (Wolfram Research, ). We provide a snippet of the Mathematica code we used below.